On January 1 2017, Headland Company sold 12% bonds having a maturity value of $ 410,000 for $. 441,084, which provides the bondholders with a 10% yield. The bonds are dated January 1, 2017, and mature January 1 2022 with interest payable December 31 of each year. Headland Company allocates interest and unamortized discount or premium on the effective-interest basis. Prepare the journal entry at the date of the bond issuance. (Round answer to O decimal places, e.g. 38,548. If no entry is required, select No Entry for the account titles and enter O for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.) Date Account Titles and Explanation Debit Credit January 1,2017 Cash 441,084 Bonds Payable 410,000 Premium on Bonds Payable 31094 eTextbook and Media List of Accounts Prepare a schedule of interest expense and bond amortization for 2017-2019. (Round answer to 0 decimal places, e.g. 38,548.) Schedule of Interest Expense and Bond Premium Amortization Effective-Interest Method Interest Expense Premium Cash Paid Carrying Amount of Bonds Date 441,084 12/31/17 12/31/18 12/31/19Prepare the journal entry to record the interest payment and the amortization for 2017. (Round answer to 0 decimal places, e.g. 38,548. If no entry is required, select No Entry for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.) Date Account Titles and Explanation Debit Credit December 31, 2017 eTextbook and Media List of Accounts Prepare the journal entry to record the interest payment and the amortization for 2019. (Round answer to O decimal places, e.g. 38,548. If no entry is required, select No Entry for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.) Date Account Titles and Explanation Debit Credit December 31,2019
Answer #1

Journal entry

Date account and explanation debit credit
Jan 1,2017 cash 441084
Bonds payable 410000
Premium on bonds payable 31084
(To record bond issue)

Schedule

Date Cash paid Interest expense Premium amortized Carrying value
1/1/2017 441084
31/12/2017 410000*12% = 49200 441084*10% = 44108 5092 435992
31/12/2018 49200 435992*10% = 43599 5601 430391
31/12/2019 49200 430391*10% = 43039 6161 424230

Journal entry

Date account and explanation debit credit
31/12/17 Interest expense 44108
Premium on bonds payable 5092
Cash 49200
(to record interest)
31/12/19 Interest expense 43039
Premium on bonds payable 6161
cash 49200
(To record interest)